OpenStudy (anonymous):

What is the yield for 6.5% interest compounded quarterly

OpenStudy (anonymous):

Know and memorize this formula for compound interest: \[A = P(1+ r/n)^{nt}\] A = Final amount P = Principal (initial) amount r = Interest rate/APR n = Times compounded per year t = Years Now having said that... There's no f-ing way I can answer your question quantitatively without you providing additional information. If you want an answer with many variables, start here, and do algebra. \[A = P(1+ 0.065/4)^{4t}\] If this is what you're having trouble with, let me know and I'd be happy to complete the rest of the problem.

OpenStudy (anonymous):

can you help me solve it please?

OpenStudy (anonymous):

Do you want the answer with all the variables? The value of A with respect to P and t?

OpenStudy (anonymous):

yes please?

OpenStudy (anonymous):

\[A = P (1+0.065/4)^{4t} = P (1.01625)^{4t} = P(1.06660161)^t\] That is your answer.. Your gross amount/yield is A.

OpenStudy (anonymous):

thank u